OpenStudy (anonymous):

Use the negative exponent rules to simplify. 9/z^-3

OpenStudy (tkhunny):

Why not use the Negative Exponent Rule? \(x^{-a} = \dfrac{1}{x^a}\) OR \(x^{a} = \dfrac{1}{x^{-a}}\)